Grade 12th passGeneral Physics

A block of mass 10 kg hanging from a political spring execute simple harmonic motion of amplitude 4 centimetre is maximum speed of political is 8 metre per second the acceleration of the block is

arham izhar
9 Years agoGrade 12th pass

1 Answer

Ritu Bisht
9 Years ago
For SHM vmax=Aw
here A=amplitude = 4cm = 0.04m
        w=angular velocity
Now  vmax=0.04*w
         8=0.04w
         w=200
Now for maximum acceleration
        amax=Aw2
               =0.04*(200)2
               =0.04*4*104
               =0.16*104
               =1600